OpenStudy (anonymous):

What is m2 if m3 = 52°? (Points : 5) 28° 52° 128° not enough information

OpenStudy (anonymous):

help please

OpenStudy (raden):

no figure , and no relation between m2 and m3 so, not enough information

OpenStudy (anonymous):

Convert to radians, calculate the result and then convert bacl to degrees. 1 radian = 180 degrees \[180\left.\left(\left(\frac{52}{180} \pi \right)^{\frac{2}{3}}\right)\right/\pi = 53.7{}^{\circ} \]